A 2300 cm3 container holds 0.12 mol of helium gas at 350 ∘C .

How much work must be done to compress the gas to 1500 cm3 at constant pressure?

How much work must be done to compress the gas to 1500 cm3 at constant temperature?
Answer #1

a)

The work done at constant pressure is given by, W = P * dV

dV = 2300 - 1500

= 800 cm3

= 0.0008 m3

By using ideal gas law:

PV = nRT

P = nRT/V

P = 0.12 * 8.314 * (623) / 0.0024

= 2..59 * 105 Pa

W = P dV

= 2..59 * 105 * 0.0008

= 207.2 Pa

b)

The work done at constant temperature is, W = n R T * ln (Vi / Vf)

W = 0.12 * 8.314 * (623) * ln (2300/1500)

= 265.68 J